# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ VI Conclusions
In general case, the magnon-photon coupling results in three magnon-polariton branches with modified dispersions. These hybrid quasiparticles exhibit remarkable properties. First of all, the magnon-polaritons carry a non-integer average spin $\langle S_{z}\rangle<\hbar$, which is tunable by both the external magnetic field and the wavevector. A notable prediction is the switching of the spin sign for the middle polariton branch upon varying the wavevector, a consequence of anticrossing mediated by the Swihart mode. The second important property is the high group velocity of the magnon-polaritons. In the strong mixing region, the group velocities of the magnon-polaritons can reach a significant fraction of the speed of light, vastly exceeding typical magnon group velocities, which promises strong tunability of magnon transport in antiferromagnets by superconductors.
@@ Acknowledgements.
Getting rid of the amplitudes $S_{1(2)y,z}$, we can rewrite Eqs. (77-80) in the following matrix form:

### v2

@@ VII Conclusions
In general case, the magnon-photon coupling results in three magnon-polariton branches with modified dispersions. These hybrid quasiparticles exhibit remarkable properties. First of all, the magnon-polaritons carry a non-integer average spin $\langle S_{z}\rangle<\hbar$, which is tunable by both the external magnetic field and the wavevector. A notable prediction is the switching of the spin sign for the middle polariton branch upon varying the wavevector, a consequence of anticrossing mediated by the Swihart mode. The second important property is the high group velocity of the magnon-polaritons. In the strong-mixing region, the group velocities of the magnon-polaritons can reach a significant fraction of the speed of light, vastly exceeding typical magnon group velocities, which promises strong tunability of magnon transport in antiferromagnets by superconductors.
@@ Acknowledgements.
Getting rid of the amplitudes $S_{1(2)y,z}$, we can rewrite Eqs. (89-92) in the following matrix form: